# Style Amnesia: Investigating Speaking Style Degradation and Mitigation in Multi-Turn Spoken Language Models

### v1

@@ 5.1 The Effect of Prompt Position
In this section, we conduct experiments under different prompt positions to investigate their effect on style amnesia. In instruction-guided language models, system messages are designed with higher priority than user messages to establish global behaviors and safety constraints Touvron et al. (2023); Wallace et al. (2024). A prior study also demonstrates that incorporating system-level instructions, such as defining personas within the system prompt, enhances multi-turn consistency and role-playing performance in LLMs Neumann et al. (2025).
Although system messages have been shown to be important for controlling LLMs, it remains unclear how the placement of speaking style instructions affects SLMs. To investigate this, we conduct an experiment comparing the performance of SLMs when instructions are placed in system messages versus user messages. We select five speaking instructions that most SLMs can follow for this experiment. The IF rates in the first turn are illustrated in Figure 4, and the full results are shown in Appendix D.5.
@@
Through the experiments above, we identify another crucial issue prevalent in current SLMs. These findings drive us to place the instruction in user messages when conducting other experiments to ensure that SLMs can follow the instruction reliably. However, results in Appendix D.5 show that style degradation occurs in both settings.
@@ 5.2 Validation of Emotion and Accent Judges
We also conduct a human validation study to assess the quality of the automatic judges for Emotion and Accent. We employ annotators via Amazon Mechanical Turk to evaluate a randomly sampled set of 720 speech clips. This subset is constructed by selecting five samples per turn across four conversational turns, six distinct styles, and six evaluated models. Each sample is evaluated by three annotators, who are asked whether the generated speech matches the required style. Details of the human evaluation setup are provided in Appendix E.1.
After collecting the annotations, we use majority vote as the final label and compute Cohen’s Kappa Cohen (1960) and Matthews Correlation Coefficient (MCC) Matthews (1975) to measure the reliability of the judges. While Cohen’s Kappa evaluates the degree of consensus among raters beyond random chance, MCC provides a balanced measure of classification quality even with imbalanced datasets. We also include Gemini-2.5 Pro Comanici et al. (2025) and VoxProfile Feng et al. (2025b) as baselines.
The results, shown in Table 3, indicate that our selected judge achieves the highest reliability. For the accent classification task, Voxlect outperforms Gemini-2.5 Pro, achieving the highest agreement with human annotations, with a Cohen’s Kappa of 0.809 and an MCC of 0.811. In the emotion classification task, Emotion2vec-Large demonstrates the strongest reliability among three emotion judges, obtaining a Cohen’s Kappa of 0.476 and an MCC of 0.511. These results align with the crowdsourced agreement levels found in common speech emotion datasets, as detailed in Appendix E.2.
@@
Figure 5: The illustration of the recall process. To save space, the instructions and conversation are shortened.
@@ 6.1 Do SLMs Forget the Instruction?
Building on the experiments in Section 4.2, we observe that SLMs suffer from style amnesia. Their speaking style following rate begins to degrade after only one turn. This raises a key research question: Do SLMs really forget the initial instructions? To address this, we introduce a recall process prior to the SLM generating speech in each turn. Specifically, for every turn after the first, we prompt the SLM to restate the initial speaking style before processing the next user input. The recall process is illustrated in Figure 5.
In the recall process, the SLM is asked to recall the original instruction using the query: “What specific instructions did the user give in the first turn regarding your speaking style for this conversation?” This allows us to verify whether the model still remember the initial instruction. We then use GPT-5 mini to judge whether the recalled instruction matches the original style instruction. The evaluation prompt is provided in Appendix C.3. We compute the percentage of cases where the SLM correctly recalls the instruction, which we define as the Recall Rate $R$.
@@
The results are summarized in Table 4. Interestingly, most SLMs remember the initial instruction quite well. The three proprietary models show a near-perfect recall rate. Qwen2.5-Omni and Step-Audio 2 mini perform worse. Their recall rates decline across turns, but still retain roughly 80% and 60% accuracy, respectively. These results reveal a clear gap between proprietary and open-source SLMs in memorization ability. From our manual inspection, when these two open-source models fail, some responses contain an incorrect style instruction, while others simply ignore the question and produce irrelevant or meaningless outputs.
@@
Notably, although GPT-4o mini exhibits a 65.3% degradation in performance when producing a sad speaking style and a 20% degradation for speaking slowly, it still maintains a high recall rate. A similar pattern is observed with Gemini Live and GPT-4o. These observations indicate that the models do not forget the instructions, suggesting that style degradation is not caused by memory loss. Instead, the models retain the instructions but fail to perform the requested style effectively.
Model Style Turn
2 3 4
Gemini Live Fast 100.0 100.0 99.0
Indian 100.0 100.0 100.0
Sadness 100.0 99.0 97.0
Slow 100.0 99.0 100.0
GPT-4o Fast 100.0 100.0 100.0
Indian 100.0 100.0 100.0
Sadness 100.0 100.0 100.0
Slow 100.0 100.0 100.0
GPT-4o mini Fast 99.0 100.0 99.0
Indian 93.0 93.0 93.8
Sadness 100.0 100.0 100.0
Slow 100.0 100.0 100.0
Qwen2.5-Omni Fast 98.0 80.0 78.0
Indian 97.0 95.0 92.0
Sadness 96.0 79.0 81.0
Slow 99.0 93.0 92.0
Step-Audio 2 mini Fast 83.0 62.0 57.0
Indian 89.0 72.0 70.0
Sadness 85.0 56.0 55.0
Slow 83.0 64.0 66.0
Table 4: Recall rate of SLMs.
6.2 Recall Prevents SLMs from Degradation
Inspired by the results from Section 6.1, we hypothesize that the degradation observed in SLMs may not stem from forgetting the initial instruction. Since our observations show that SLMs actually retain the instruction when prompted, we investigate whether explicitly asking them to recall the instruction can mitigate style degradation. We conduct the following experiments to examine whether the recall process can mitigate style amnesia issues. We conduct experiments on models capable of producing a wide range of speaking styles, as well as on styles that most models can reliably generate.
@@ 6.2 Recall Prevents SLMs from Degradation
The results, presented in Table 5, show that SLMs equipped with the recall process notably reduce degradation. Even for models with relatively low degradation, such as Gemini Live and GPT-4o, the recall process still provides measurable improvements. GPT-4o mini, which suffers from substantial degradation, achieves roughly a 25% reduction in the average degradation rate, demonstrating the effectiveness of the recall process. Step-Audio 2 mini shows slightly improved but largely comparable degradation across the four tasks, likely due to its lower recall rate relative to other SLMs.
@@
Although the recall process alleviates style degradation, some degree of decay remains unavoidable. This suggests that the inconsistency originates from fundamental limitations of current SLMs and must be addressed in future model designs. As shown in Figure 1 and in Appendix D.1, degradation emerges after just a single turn. This early degradation explains why the recall process cannot fully resolve the problem. It still depends heavily on the SLM’s inherent ability to maintain style consistency.
Model Sadness Indian Fast Slow Avg
Gemini Live 21.3 0.0 9.7 0.7 7.9
   + Recall 17.3 0.0 6.3 0.7 6.1
   Improvement +4.0 0.0 +3.4 0.0 +1.8
GPT-4o 26.7 1.7 3.3 8.1 10.0
   + Recall 14.9 0.0 0.3 4.4 4.9
   Improvement +11.8 +1.7 +3.0 +3.7 +5.1
GPT-4o mini 65.3 49.7 4.0 20.9 35.0
   + Recall 30.3 14.9 0.0 1.5 11.7
   Improvement +35.0 +34.8 +4.0 +19.4 +23.3
Step-Audio 2 mini 14.0 5.0 27.0 27.0 18.3
   + Recall 11.0 5.3 29.0 24.7 17.5
   Improvement +3.0 -0.3 -2.0 +2.3 +0.8
@@
Table 5: Impact of Recall Process on Degradation Rate $D$ across different styles.
@@ 7 Conclusion
In this paper, we identify that SLMs are prone to getting lost in multi-turn conversations, which leads to inconsistent speaking styles across turns. We also demonstrate that SLMs may remember the instruction when asked, but fail to perform it explicitly. Additionally, we demonstrate that placing style instructions in system messages results in a significant performance drop.
Built upon a robust text-only LLM backbone, their ability to memorize speaking style is already relatively strong. However, since the text-only LLM lacks the module to adjust speaking style, SLMs’ ability to maintain consistent style gains little benefit from text-only pretraining. These observations reflect that further learning is required when developing SLMs. We hope these findings offer valuable insights for the community and contribute to the development of more reliable SLMs.
@@
Nonetheless, these limitations do not weaken the conclusions of our study. Even with the attributes we can evaluate, SLMs already exhibit noticeable degradation in multi-turn dialogues, suggesting that similar or even greater challenges may arise in more complex settings. We therefore leave the analysis of these scenarios to future work, as it requires further advances in both SLMs and judge models.
@@ Appendix B Dataset Construction
To minimize topic-induced variance by averaging across diverse conversation contents, we select samples from the SODA dataset Kim et al. (2023) to generate conversation openers. SODA is an English dialogue dataset covering a wide range of social interactions. Each dialogue in SODA includes a narrative context, speaker name, and a knowledge graph that defines the events and relationships within the dialogue.
@@ D.1 Full Main Results
The IF rate for each turn are shown in Table 7, showing that all evaluated SLMs exhibit style degradation.
@@ D.2 SLM Default Speaking Styles
We observe that most SLMs show higher consistency when generating Happiness, Neutral tone, and North American English than other style attributes. This is likely because these attributes match the models’ default speaking styles. To verify this, we examine the emotion and accent distributions of samples generated during the speed and volume evaluations. As shown in Table 6, most SLMs tend to perform well with happy or neutral tones and North American accents, which explains the less degradation for these styles.
@@ D.3 Dialogue Coherence Evaluation
D.4 Prompt Paraphrasing
To further investigate how the phrasing of instructions affects the results, we conduct an ablation study on Step-Audio 2 mini. The original instruction template we use in other experiments is “In the following conversation, you must consistently…” We prepare two variant instruction templates. The first variant simply rewrites the instruction to be “From now on, you can only…”. The second variant uses a persona-based template, "You are a person who always…, no matter what." and instructs SLMs to speak at a fast and slow pace, respectively.
The results about speaking rate are illustrated in Figure 6. In Figure 6, we plot the mean and the variance of WPM for each turn. As the figure shows, there is nearly no difference between the different instructions. This experiment demonstrates that the style degradation is not caused by the prompt paraphrasing.
Figure 6: The results of different instructions. The line represents the mean across all samples, and the shaded background area indicates the range of the mean $\pm$ standard deviation. Instruction 1 is the prompt used in the other experiments.
D.5 Prompt Position
@@
D.6 GPT-4o Fails to Return Speech
@@
Lin et al. (2025) indicates that GPT-4o exhibited a high refusal rate. We hypothesize that the issue stems from the model’s internal safety guard mechanisms. Based on our observations, it is possible that these two models return no speech at each round. Since they are proprietary models to which we do not have access, it is difficult for us to definitively address this issue. As the percentage of failed cases is low and the observed style degradation is significant, we believe this issue will not change our conclusions.
@@
Appendix E Automatic Judge Validation
E.1 Human Evaluation Setups
  • Annotators must have an approval rate higher than 98%.
  • Annotators must have more than 10,000 approved tasks.
@@ E.2 Inter-Annotator Agreement of Speech Emotion Recognition Datasets
IEMOCAP Busso et al. (2008), CREMA-D Cao et al. (2014), and MELD Poria et al. (2019) are three widely used speech emotion recognition datasets. These datasets provide large-scale speech emotion annotations through crowdsourcing, making significant contributions to the development of speech emotion recognition models. Emotion in speech is inherently subjective and may be perceived differently by different annotators. As a result, inter-annotator agreement (IAA) is often imperfect. This variability does not indicate annotation noise or low quality, but reflects the inherent subjectivity of emotion perception in speech.
Among them, IEMOCAP and MELD adopt Fleiss’ Kappa Fleiss (1971) as the IAA indicator. Fleiss’ Kappa is designed to assess the reliability of agreement among a fixed number of annotators assigning categorical labels to a set of items, measuring the extent to which the observed agreement exceeds what would be expected by chance. Each sample in the IEMOCAP dataset is annotated by three evaluators, yielding a Fleiss’ Kappa of 0.27, whereas the MELD dataset demonstrates a higher agreement of 0.43 among three annotators.
In comparison, the CREMA-D dataset uses Krippendorff’s alpha Krippendorff (1980) to report the agreement. This metric is used to measure agreement across any number of observers and is particularly robust because it can handle incomplete data and various data scales while accounting for disagreements expected by chance. The authors collect samples with an average of 9.8 annotators per clip and exhibit a Krippendorff’s alpha of 0.42. Furthermore, they report a self-consistency rate of approximately 70%, which represents the level of agreement when the same annotator evaluates the same sample at different times. In sum, these scores highlight the inherent subjectivity and complexity of emotional annotation in spoken dialogue.

### v2

@@ 5.1 The Effect of Prompt Position
In this section, we conduct experiments under different prompt positions to investigate their effect on style amnesia. In instruction-guided language models, system messages are designed with higher priority than user messages to establish global behaviors and safety constraints touvron2023llama2; wallace2024instruction. A prior study also demonstrates that system prompts have a more profound impact on model behavior than user prompts do. 10.1145/3715275.3732038.
Although system messages are important for controlling LLMs, it remains unclear how the placement of speaking style instructions affects SLMs. To investigate this, we conduct an experiment comparing the performance of SLMs when instructions are placed in system messages versus user messages. We select five speaking instructions that most SLMs can follow for this experiment. The IF rate in the first turn is illustrated in Figure 5, and the IF rate for each turn is shown in Appendix D.4.
@@
The results presented here focus on the IF rate in the first turn. However, the IF rate for each turn reported in Appendix D.4 show that style amnesia occurs in both settings. Through the experiments above, we identify another crucial issue prevalent in current SLMs. These findings drive us to place the instruction in user messages when conducting other experiments, ensuring that SLMs can follow the instructions more effectively.
@@ 5.2 Validation of Emotion and Accent Judges
As described in Section 3.3.1, we evaluate Speed and Volume using deterministic, signal-level metrics. In contrast, Emotion and Accent are evaluated using learned models designed to approximate human perception. To ensure the reliability of the automatic judges for Emotion and Accent, we conduct a human validation study to assess the correlation between our automatic judges and human annotators. In addition, we compare our automatic judges with Gemini-2.5 Pro comanici2025gemini, a robust LALM capable of general-purpose speech understanding. This model has been shown to exhibit high correlation with human annotators when used as an automatic judge in a prior LALM-as-a-judge study chiang2025audio.
We employ annotators via Amazon Mechanical Turk to evaluate a randomly sampled subset of 720 speech clips. This subset is constructed by selecting five samples per turn across four conversational turns, six speaking styles, and six evaluated models. The six evaluated styles include Happiness, Neutral, Anger, Sadness, Indian English accent, and North American English accent. Each sample is evaluated by three annotators, who are asked whether the generated speech matches the required style. Details of the human evaluation setup are provided in Appendix E.1.
@@
After collecting the annotations, we derive the final label using majority voting, compute Cohen’s Kappa cohen1960coefficient for inter-annotator agreement, and use Matthews Correlation Coefficient (MCC) matthews1975comparison to evaluate judge reliability against the final labels. While Cohen’s Kappa evaluates the degree of consensus among raters, MCC provides a balanced measure of classification quality even with imbalanced datasets. Therefore, we report both metrics for reference.
The results, shown in Table 2, indicate that our selected judge achieves the highest reliability. For the accent classification task, Voxlect outperforms Gemini-2.5 Pro, achieving the highest agreement with human annotations, with a Cohen’s Kappa of 0.809 and an MCC of 0.811. In the emotion classification task, Emotion2vec-Large demonstrates the strongest reliability among three emotion judges, obtaining a Cohen’s Kappa of 0.476 and an MCC of 0.511. These results are similar to the crowdsourced agreement levels found in common speech emotion datasets, as detailed in Appendix E.2.
Figure 6: The illustration of the recall process.
@@ 6 SLM Recall Process
Building on the experiments in Section 4.2, we observe that SLMs suffer from style amnesia: their speaking style IF rate begins to degrade after only one turn and often worsens as the conversation progresses. This raises a key research question: Do SLMs forget the initial instruction, or just fail to follow the specified style? To address this, we introduce a recall process for every turn after the first, in which the SLM is prompted to restate the initial speaking style before processing the following user input. The recall process is illustrated in Figure 6. Using this method, we first utilize the recall probe to measure whether the SLM retains the original instruction in Section 6.1. Subsequently, we evaluate the effectiveness of integrating this recall process to mitigate style amnesia in Section 6.2.
@@ 6.1 Do SLMs Forget the Instruction?
To quantify whether the model remembers the initial style instruction $s$, we define the recall rate $R$ as follows. Given a style $s$ and dialogue topic $i$, the recall rate at assistant turn $j$, denoted $R_{j}(s)$, is:
$R_{j}(s)=\frac{\sum_{i=1}^{N}\mathbbm{1}_{\text{recall}}(r_{i,j},s)}{N}\times 100\%,$ (3)
where $r_{i,j}$ denotes the response generated by the SLM to the recall query $q_{\text{recall}}$ before generating its response at the user turn $j$, and $N$ is the number of dialogue topics. The binary indicator function $\mathbbm{1}_{\text{recall}}(\cdot)$ is 1 if the recalled instruction matches the original style instruction, and 0 otherwise. We use GPT-5 mini to judge the correctness of the recalled instruction. The evaluation prompt is shown in Appendix C.3.
@@
We conduct the following experiments on models capable of producing a wide range of speaking styles, as well as on styles that most models can reliably generate. The results are summarized in Table 3. Interestingly, most SLMs remember the initial instruction quite well. The three proprietary models show a near-perfect recall rate. Step-Audio 2 mini shows weaker performance with declining recall rates across turns, yet it still achieves a recall rate of 55.0% to 89.0%. These results reveal a clear gap between proprietary and open-source SLMs in memorization ability. From our manual inspection, when Step-Audio 2 mini fails to recall, some responses contain an incorrect style instruction, while others simply ignore the question and produce irrelevant or meaningless outputs.
@@
Notably, although GPT-4o mini exhibits a 65.3% degradation in performance when producing a sad speaking style and a 20% degradation for speaking slowly, it still maintains a high recall rate. A similar pattern is observed with Gemini Live and GPT-4o. These observations indicate that the models do not forget the instructions, suggesting that style amnesia is not caused by memory loss. Instead, the models retain the instructions but fail to follow the requested style effectively.
6.2 Recall Process Mitigates Style Amnesia
@@
In this section, we investigate whether explicitly asking SLMs to recall the instruction can mitigate style amnesia. The results, presented in Table 3, show that SLMs equipped with the recall process notably reduce degradation. Even for models with relatively low degradation, such as Gemini Live and GPT-4o, the recall process still provides measurable improvements. GPT-4o mini, which suffers from substantial degradation, achieves roughly a 25% reduction in the average degradation rate, demonstrating the effectiveness of the recall process. Step-Audio 2 mini shows slightly improved but largely comparable degradation across the four tasks, likely due to its lower recall rate relative to other SLMs.
@@
Although the recall process alleviates style amnesia, some degree of decay remains unavoidable. This suggests that the inconsistency originates from fundamental limitations of current SLMs and must be addressed in future model designs.
Model Style $R_{2}$ $R_{3}$ $R_{4}$ Degradation Rate $D$
Base +Recall Improv.
Gemini Live Indian 100.0 100.0 100.0 0.0 0.0 0.0
Sadness 100.0 99.0 97.0 21.3 17.3 +4.0
Fast 100.0 100.0 99.0 9.7 6.3 +3.4
Slow 100.0 99.0 100.0 0.7 0.7 0.0
GPT-4o Indian 100.0 100.0 100.0 1.7 0.0 +1.7
Sadness 100.0 100.0 100.0 26.7 14.9 +11.8
Fast 100.0 100.0 100.0 3.3 0.3 +3.0
Slow 100.0 100.0 100.0 8.1 4.4 +3.7
GPT-4o mini Indian 93.0 93.0 93.8 49.7 14.9 +34.8
Sadness 100.0 100.0 100.0 65.3 30.3 +35.0
Fast 99.0 100.0 99.0 4.0 0.0 +4.0
Slow 100.0 100.0 100.0 20.9 1.5 +19.4
Step-Audio 2 mini Indian 89.0 72.0 70.0 5.0 5.3 -0.3
Sadness 85.0 56.0 55.0 14.0 11.0 +3.0
Fast 83.0 62.0 57.0 27.0 29.0 -2.0
Slow 83.0 64.0 66.0 27.0 24.7 +2.3
@@
Table 3: The recall rate $R$ of SLMs and the impact of the recall process on the degradation rate $D$ across different styles. Base represents direct inference without the recall process, while + Recall indicates performance using the SLM integrated with the recall process.
@@ 7 Conclusion
In this paper, we identify that SLMs suffer from style amnesia in multi-turn conversations, which leads to inconsistent speaking styles across turns. We also demonstrate that SLMs can recall the instruction when asked, but fail to perform it explicitly. Additionally, we indicate that placing style instructions in system messages results in a significant performance drop.
SLMs exhibit a relatively strong ability to memorize specified speaking styles. However, the retention of stylistic information does not necessarily translate into stylistic expression at generation time. Closing this gap between style retention and stylistic control remains an important direction for future research. We hope these findings offer valuable insights for the community and contribute to the development of more reliable SLMs.
@@
Nonetheless, these limitations do not weaken the conclusions of our study. Even with the styles we can evaluate, SLMs already exhibit noticeable degradation in multi-turn dialogues, suggesting that similar or even greater challenges may arise in more complex settings. We therefore leave the analysis of these scenarios to future work, as it requires further advances in both SLMs and judge models.
We do not see specific harm in our paper.
@@ Appendix B Dataset Construction
To minimize topic-induced variance by averaging across diverse conversation contents, we select samples from the SODA dataset kim2023soda to generate conversation openers. Soda is an English dialogue dataset covering a wide range of social interactions. Each dialogue in SODA includes a narrative context, speaker name, and a knowledge graph that defines the events and relationships within the dialogue. Soda is released under the Creative Commons Attribution 4.0 (CC BY 4.0) license. Accordingly, our use of the dataset is compliant with the license requirements, including proper attribution to the original authors.
@@ D.1 Full Main Results
The IF rate for each turn are shown in Table 5, showing that all evaluated SLMs exhibit style amnesia.
@@ D.2 SLM Default Speaking Styles
We observe that most SLMs show higher consistency when generating Happiness, Neutral tone, and North American English than with other style attributes. This is likely because these attributes match the models’ default speaking styles. To verify this, we examine the emotion and accent distributions of samples generated during the speed and volume evaluations. As shown in Table 4, most SLMs tend to perform well with happy or neutral tones and North American accents, which explains the less degradation for these styles.
@@ D.3 Dialogue Coherence Evaluation
D.4 Prompt Position
@@
D.5 GPT-4o Fails to Return Speech
@@
lin2025preliminary indicate that GPT-4o exhibited a high refusal rate. We hypothesize that the issue stems from the model’s internal safety guard mechanisms. Based on our observations, it is possible that these two models return no speech at each round. Since they are proprietary models to which we do not have access, it is difficult for us to definitively address this issue. As the percentage of failed cases is low and the observed style degradation is significant, we believe this issue will not change our conclusions.
Appendix E Automatic Judge Validation
E.1 Human Evaluation Setups
  • Annotators must have an approval rate higher than 98%.
  • Annotators must have more than 10,000 approved tasks.
IEMOCAP busso2008iemocap, CREMA-D cao2014crema, and MELD poria2019meld are three widely used speech emotion recognition datasets. These datasets provide large-scale speech emotion annotations through crowdsourcing, making significant contributions to the development of speech emotion recognition models. Emotion in speech is inherently subjective and may be perceived differently by different annotators. As a result, inter-annotator agreement (IAA) is often imperfect. This variability does not indicate annotation noise or low quality, but reflects the inherent subjectivity of emotion perception in speech.
Among them, IEMOCAP and MELD adopt Fleiss’ Kappa fleiss1971measuring as the IAA indicator. Fleiss’ Kappa is designed to assess the reliability of agreement among a fixed number of annotators assigning categorical labels to a set of items, measuring the extent to which the observed agreement exceeds what would be expected by chance. Each sample in the IEMOCAP dataset is annotated by three evaluators, yielding a Fleiss’ Kappa of 0.27, whereas the MELD dataset demonstrates a higher agreement of 0.43 among three annotators.
In comparison, the CREMA-D dataset uses Krippendorff’s alpha krippendorff2018content to report the agreement. This metric is used to measure agreement across any number of observers and is particularly robust because it can handle incomplete data and various data scales while accounting for disagreements expected by chance. The authors collect samples with an average of 9.8 annotators per clip and exhibit a Krippendorff’s alpha of 0.42. Furthermore, they report a self-consistency rate of approximately 70%, which represents the level of agreement when the same annotator evaluates the same sample at different times. In sum, these scores highlight the inherent subjectivity and complexity of emotional annotation in spoken dialogue.